# Remote Cooling of Spin-ensembles through a Spin-mechanical Hybrid Interface

### v2

@@ IV System cooling
In this section, we describe how the repeated probe projections can impose a thermal filter on the hybrid system, allowing for controlled cooling of either the oscillator, the ensemble, or both. We focus primarily on the weak coupling regime, where the coupling strength $g$ is significantly smaller than the mechanical oscillation frequency $\omega$, denoted as $g\ll\omega$. To explore this regime, we set the spin coupling strengths $g,~{}g_{0}$ to $10$ kHz, while the mechanical oscillation frequency $\omega$ is chosen to be $1.2$ MHz. Although such a frequency for the mechanical oscillator is experimentally achievable [45, 46], the chosen values for the spin-mechanical couplings are at the upper end of what is typically achievable in practice [43].
These choices represent a compromise between stronger coupling for faster cooling and additional noise due to imperfect pulses on the probe. As we discussed in Eq. (11), at least $\mathcal{O}(\omega/g_{0})$ pulses are required for each probe projection to ensure that the effective spin-oscillator coupling is comparable to the oscillator frequency, which becomes clear in Eq. (9). However, a higher number of pulses would result in a longer total evolution time to implement the cooling protocol, which is undesirable in the presence of decoherence. Throughout this work, we set the number of pulses in each round of probe projection to $n=\omega/g_{0}=120$, as shown in Fig. 1(b).
@@
Figure 3: Pulse Detuning: A key parameter in the cooling scheme. (a) This panel shows the intricate dependence of achieving complete spin polarization on pulse detuning ($\epsilon$). The probability that all ensemble spins are aligned upward after $M=100$ probe projections, denoted by $\mathrm{P}_{\scriptscriptstyle m=N/2}(M)$ and derived in Eq. (16), is plotted on the vertical axis. This probability is affected by both the number of spins in the ensemble and the size of $\epsilon$, with $\epsilon=2$ kHz already appearing too large for effective polarization of the ensemble. (b) If we choose an even larger detuning ($\epsilon=2.5$ kHz) and set the number of spins to $N=4$, we observe a rapid decrease of the thermal occupancy of the oscillator to zero, indicating that the oscillator is cooling. However, the spin polarization remains almost unchanged and reaches a steady state of 0.0625 after about 40 probe projections. Note that we have considered an ensemble of $N=4$ spins, so the y-axis represents the relative value of $\mathrm{P}_{z}(M)/2$. Correspondingly, the y-axis also represents the relative value of the thermal occupancy of the oscillator with respect to its initial value, which is set to about 45. Further simulation details are summarized in Appendix A.
@@ A Pulse detuning as a control parameter
First, we run simulations with the pulse detuning $\epsilon$ close to zero to explore the near-resonance regime. By adjusting its magnitude from 0 to 2 kHz and varying the number of spins in the ensemble, our simulations reveal a nuanced relationship between the probability of achieving complete ensemble polarization and the pulse detuning $\epsilon$, as shown in Fig. 3(a). To cool the spin ensemble, the choice of $\epsilon$ should satisfy the condition that the probe-oscillator coupling is enhanced without prematurely cooling the oscillator. Consequently, satisfying this condition requires fine-tuning of all parameters of the Hamiltonian. This makes it difficult to determine the exact dependence of the optimal pulse detuning $\epsilon$ on the number of ensemble spins $N$ and the number of probe projections $M$, as shown in Fig. 3(a). For example, it can be seen that within an ensemble of constant size, varying the detuning can lead to very different results when implementing the proposed cooling protocol. Another important observation from Fig. 3(a) suggests that a detuning of $\epsilon=2$ kHz may be large enough to facilitate rapid oscillator cooling and thereby prevent ensemble polarization.
To investigate this further, we introduce an even larger detuning of $\epsilon=2.5$ kHz for an ensemble of $N=4$ spins. The simulation results shown in Fig. 3(b) indicate that at this substantial pulse detuning, the oscillator quickly reaches its ground state after about 80 probe projections. This premature cooling of the oscillator reduces the interaction between the probe and the ensemble, so that the ensemble remains largely unpolarized. The ensemble polarization fluctuates only slightly around the final saturation value for the first few tens of probe projections, as shown in the inset of Fig. 3(b).
@@
This observation is interestingly related to the encoding of Gottesman-Kitaev-Preskill (GKP) states, which have applications in quantum sensing and quantum error correction [47, 48, 49]. Its encoding can be realized by the same control of the probe spin as our cooling protocol. Starting with an oscillator in its zero photon ground state, repeated probe projections can actually increase the number of photons and leave the oscillator in GKP states [47, 48]. The generation of these states in macroscopic objects could enrich our understanding of the quantum-classical interface, notwithstanding their important roles in quantum error correction [47, 48] and quantum sensing [50]. More details of GKP encoding can be found in Appendix C.
Figure 4: Cooling of both the oscillator and the ensemble. This figure shows the polarization $\mathrm{P}_{z}(M)$ and the thermal occupancy $n(\omega)$ for both the spin ensemble and the oscillator, illustrating how the cooling/polarization proceeds with the number of probe projections $M$. The effective cooling of these quantum systems occurs in different parameter regimes. Initially, the spin ensemble is cooled by adjusting the pulse intervals to near-resonance conditions, specifically $\tau_{1}$ such that $\epsilon_{1}=1.1$ kHz. In this case, the thermal occupancy of the oscillator remains largely unaffected. After polarizing the spin ensemble, we change the pulse interval to $\tau_{2}$ with a pulse detuning of $\epsilon_{2}=5$ kHz. This more significant detuning leads to a rapid cooling of the oscillator to its ground state. Note that the number of ensemble spins is $N=4$ so that the y-axis represents the relative value of spin polarization $\mathrm{P}_{z}(M)/2$. The oscillator thermal occupancy is also plotted as its relative to the initial occupancy $n_{0}\approx 45$. Further simulation details are summarized in Appendix A.
B Simultaneous cooling of ensemble and oscillator
@@
During this simulation, we find that the optimal detuning that achieves complete polarization of the ensemble is identified as $\epsilon_{1}=1.1$kHz, while the oscillator occupancy is not much reduced. In this case, performing 250 probe projections can achieve complete polarization of the ensemble. Then, to cool the oscillator, we adjust the detuning to a higher value, $\epsilon_{2}=5$kHz, and execute an additional series of $50$ probe projections. This transition in the magnitude of the pulse detuning results in a rapid cooling of the oscillator, as shown in the right part of Fig. 4.
@@ V Discussion
In this study, we have developed a novel ground-state cooling protocol for a tripartite hybrid quantum system consisting of a spin ensemble, an oscillator, and a probe spin. Central to this protocol is the exploitation of the feedback provided by frequent measurements of the prob spin. Our analysis has shown that even with weak dispersive coupling between the spins and the oscillator, one can effectively cool both the ensemble and the oscillator to their ground states, either separately or simultaneously. We would also like to emphasize that, except for the probe, there is no need for quantum coherence in other parts of the hybrid system, marking a clear distinction from resonant transfer schemes. The polarization considered here can also be viewed as a remote sensing scheme when considering a macroscopic oscillator, where the boundaries of the sensing volume can be extended beyond the limits set by phase accumulation schemes [25].
@@
Our numerical simulations, performed over a range of parameters, underscore the feasibility of the cooling protocol, and lay the groundwork for its experimental realization. For example, our results indicate that a spin ensemble consisting of $N=14$ spins could be cooled with a spin-mechanical coupling strength of only a few kHz at a mechanical oscillation frequency close to 1 MHz. Although such coupling strengths are within the experimental reach, the cooling efficiency can be further improved by miniaturization of the mechanical oscillator [51]. In the context of solid-state defects in silicon carbide, coupling strengths can reach a few MHz [37], promising even more efficient cooling rates. In addition, the use of magnons instead of NMOs could significantly enhance the spin-oscillator coupling, facilitating the experimental adoption of our protocol.
@@
A notable challenge of our cooling approach is its dependence on post-selection, which results in a decreasing success rate with increasing number of probe projections. Nevertheless, our protocol offers a significant advantage for cooling large spin ensembles that are traditionally difficult to handle and cannot be cooled by standard optical or energy transfer techniques. For spin ensembles that are amenable to conventional cooling methods, such as NV centers in diamond, our protocol adds to the cooling toolbox.
@@
This research highlights the significant impact that precise control of a single spin can have on the dynamics of a hybrid system, enabling energy transfer-like effects beyond the inherent capabilities of weak dispersive coupling. In addition to remote sensing, these results also provide insight into the preparation of complex quantum states such as the GKP states in a macroscopic object. Furthermore, the ability to interact/communicate with a few target spins within a large network may also be useful in quantum network protocols where a similar result is desired.
@@ Acknowledgement
The authors would like to thank Jixing Zhang, Xin Zhang, and Ruoming Peng for their valuable feedback on the manuscript. This work was supported by the German Federal Ministry of Education and Research (BMBF) through the project QECHQS with Grant No. 16KIS1590K, and through the Future Cluster QSens. Additional support was provided by the Baden-Württemberg Stiftung through the project SPOC with Grant agreement No. QT-6. The Deutsche Forschungsgemeinschaft (DFG) has funded this work under GRK 2642, and under Project No. 384846402 as part of FOR 2724.
Appendix A Simulation details

### v3

@@ IV System cooling
In this section, we demonstrate the ability of repeated probe projections to impose a thermal filter on the hybrid system, allowing controlled cooling of either the oscillator, the ensemble, or both. In the case of simultaneous cooling, the oscillator can only be cooled after the ensemble is fully polarized; otherwise, there is effectively no interaction between the probe and the ensemble. As illustrated in Fig. 4, the cooling process for the entire system is divided into two stages with distinct pulse spacings $\tau$. In the first stage, the oscillator remains in its thermal state while the ensemble is cooled. In the second stage, another pulse spacing is used to rapidly cool the oscillator, following the oscillator cooling scheme described in our previous work [28].
We explore a range of parameters to identify this effect through numerical simulations, focusing primarily on the weak coupling regime where the coupling strength $g$ is significantly smaller than the mechanical oscillation frequency $\omega$, denoted as $g\ll\omega$. Details about these simulations are summarized in Appendix C.
@@
To explore this weak-coupling regime, we set the spin coupling strengths $g,~{}g_{0}$ to $10$ kHz, while the mechanical oscillation frequency $\omega$ is chosen to be $1.2$ MHz. Although such a frequency for the mechanical oscillator is experimentally achievable [33, 34], the chosen values for the spin-mechanical couplings are at the upper end of what is typically achievable in practice [28].
@@
Figure 3: Pulse Detuning: A key parameter in the cooling scheme. (a) This panel shows the intricate dependence of achieving complete spin polarization on pulse detuning ($\epsilon$). The probability that all ensemble spins are aligned upward after $M=100$ probe projections, denoted by $\mathrm{P}_{\scriptscriptstyle m=N/2}(M)$ and derived in Eq. (14), is plotted on the vertical axis. This probability is affected by both the number of spins in the ensemble and the size of $\epsilon$, with $\epsilon=2$ kHz already appearing too large for effective polarization of the ensemble. (b) If we choose an even larger detuning ($\epsilon=2.5$ kHz) and set the number of spins to $N=4$, we observe a rapid decrease of the thermal occupancy of the oscillator to zero, indicating that the oscillator is cooling. However, the spin polarization remains almost unchanged and reaches a steady state of 0.0625 after about 40 probe projections. Note that we have considered an ensemble of $N=4$ spins, so the y-axis represents the relative value of $\mathrm{P}_{z}(M)/2$. Correspondingly, the y-axis also represents the relative value of the thermal occupancy of the oscillator with respect to its initial value, which is set to about 45. Further simulation details are summarized in Appendix C.
These choices represent a compromise between stronger coupling for faster cooling and additional noise due to imperfect pulses on the probe. As we discussed in Eq. (11), at least $\mathcal{O}(\omega/g_{0})$ pulses are required for each probe projection to ensure that the effective spin-oscillator coupling is comparable to the oscillator frequency, which becomes clear in Eq. (9). However, a higher number of pulses would result in a longer total evolution time to implement the cooling protocol, which is undesirable in the presence of decoherence. Throughout this work, we set the number of pulses in each round of probe projection to $n=\omega/g_{0}=120$, ensuring that the effective probe-oscillator coupling strength is sufficiently strong to drive the oscillator, as demonstrated in Eq. (9).
In this work, we refrain from exhaustively exploring all possible parameter regimes. Instead, we focus on fixed coupling strengths and oscillator frequencies that are experimentally feasible, and consider only a few spins in the ensemble. Our primary goal is to demonstrate that manipulating the probe spin, specifically by tuning the pulse detuning amplitude $\epsilon$, can significantly alter the dynamics of a hybrid oscillator-ensemble system with only weak dispersive couplings. The parameter regimes we explore are sufficient to confirm this capability, as we will discuss below.
@@ A Pulse detuning as a control parameter
First, we run simulations with the pulse detuning $\epsilon$ close to zero to explore the near-resonance regime. By adjusting its magnitude from 0 to 2 kHz and varying the number of spins in the ensemble, our simulations reveal a nuanced relationship between the probability of achieving complete ensemble polarization and the pulse detuning $\epsilon$, as shown in Fig. 3(a). To cool the spin ensemble, the choice of $\epsilon$ should satisfy the condition that the probe-oscillator coupling is enhanced without prematurely cooling the oscillator. Consequently, satisfying this condition requires fine-tuning of all parameters of the Hamiltonian.
This makes it difficult to determine the exact dependence of the optimal pulse detuning $\epsilon$ on the number of ensemble spins $N$ and the number of probe projections $M$, as shown in Fig. 3(a). For example, it can be seen that within an ensemble of constant size, varying the detuning can lead to very different results when implementing the proposed cooling protocol. Another important observation from Fig. 3(a) suggests that a detuning of $\epsilon=2$ kHz may be large enough to facilitate rapid oscillator cooling and thereby prevent ensemble polarization.
B Premature oscillator cooling
To investigate the premature cooling of the oscillator further, we introduce an even larger detuning of $\epsilon=2.5$ kHz for an ensemble of $N=4$ spins. The simulation results shown in Fig. 3(b) indicate that at this substantial pulse detuning, the oscillator quickly reaches its ground state after about 80 probe projections. This premature cooling of the oscillator reduces the interaction between the probe and the ensemble, so that the ensemble remains largely unpolarized. The ensemble polarization fluctuates only slightly around the final saturation value for the first few tens of probe projections, as shown in the inset of Fig. 3(b).
@@ B Premature oscillator cooling
This observation is also interestingly related to the encoding of Gottesman-Kitaev-Preskill (GKP) states, which have applications in quantum sensing and quantum error correction [35, 36, 37]. Its encoding can be realized by the same control of the probe spin as our cooling protocol. Starting with an oscillator in its zero photon ground state, repeated probe projections can actually increase the number of photons and leave the oscillator in GKP states [35, 36]. The generation of these states in macroscopic objects could enrich our understanding of the quantum-classical interface, notwithstanding their important roles in quantum error correction [35, 36] and quantum sensing [38]. More details of GKP encoding can be found in Appendix D.
Figure 4: Cooling of both the oscillator and the ensemble. This figure shows the polarization $\mathrm{P}_{z}(M)$ and the thermal occupancy $n(\omega)$ for both the spin ensemble and the oscillator, illustrating how the cooling/polarization proceeds with the number of probe projections $M$. The effective cooling of these quantum systems occurs in different parameter regimes. Initially, the spin ensemble is cooled by adjusting the pulse intervals to near-resonance conditions, specifically $\tau_{1}$ such that $\epsilon_{1}=1.1$ kHz. In this case, the thermal occupancy of the oscillator remains largely unaffected. After polarizing the spin ensemble, we change the pulse interval to $\tau_{2}$ with a pulse detuning of $\epsilon_{2}=5$ kHz. This more significant detuning leads to a rapid cooling of the oscillator to its ground state. Note that the number of ensemble spins is $N=4$ so that the y-axis represents the relative value of spin polarization $\mathrm{P}_{z}(M)/2$. The oscillator thermal occupancy is also plotted as its relative to the initial occupancy $n_{0}\approx 45$. Further simulation details are summarized in Appendix C.
C Simultaneous cooling of ensemble and oscillator
@@
During this simulation, we find that the optimal detuning that achieves complete polarization of the ensemble is identified as $\epsilon_{1}=1.1$kHz, while the oscillator occupancy is not much reduced. In this case, performing 200 probe projections can achieve complete polarization of the ensemble. Then, to cool the oscillator, we adjust the detuning to a higher value, $\epsilon_{2}=5$kHz, and execute an additional series of $50$ probe projections. This transition in the magnitude of the pulse detuning results in a rapid cooling of the oscillator, as shown in the right part of Fig. 4.
@@
D Effects of decoherence
Our numerical simulations have not considered the direct effect of spin dephasing noise. For the hybrid system considered in this work, such noise is induced by the spin bath surrounding the probe and the ensemble and is typically the dominant noise source [26, 39]. Firstly, the repeated inversion pulses effectively protect the probe spin from dephasing noise. Notably, this technique can extend the coherence time of an NV electronic spin to exceed one second under cryogenic temperatures [39]. Consequently, such a long coherence time is not a limiting factor for our protocol. Additionally, after each probe projection, both the spin ensemble and its surrounding spin bath gradually collapse towards their steady states simultaneously, a process that is sensitive only to the probe population in its measurement basis [40]. One thing worth noting is that previous works, such as Refs. [41, 22], have counterintuitively shown that decoherence of the spins in an ensemble effectively enhances its polarization rather than decreasing it.
Similarly, regarding the oscillator cooling, the effects of the oscillator dissipation ($T_{1}$) is also negligible. As shown in Fig. 4, the initial bosonic occupancy of the oscillator remains almost unaffected during ensemble cooling. Therefore, the effect of the oscillator’s $T_{1}$ on ensemble cooling can be ignored. Once the ensemble is fully polarized, there is only one spin configuration with all spins in the same state. The spin ensemble now exerts only a static force on the oscillator, displacing the oscillator to a new equilibrium. The right panel of Fig. 4 hence exactly corresponds to the oscillator cooling scheme presented in our previous work [28], where it was shown that the $(T_{1})$ decay of the oscillator only influences the total number of probe projections required to achieve a similar cooling effect i.e., $T_{1}$ decreases the cooling rate but not the achievable cooling limit.
We note that other imperfections, such as inhomogeneous ensemble-oscillator coupling and imperfect probe readout, will affect the cooling efficiency but not the final state that can be achieved. A more detailed discussion can be found in the Appendix.
E Cooling efficiency
It is crucial to note that our protocol only accepts a probe measurement result of 0; any other outcome necessitates restarting the entire process. This requirement leads to an exponentially decreasing success rate, which constitutes the primary limitation in achieving full polarization of a spin ensemble in practice. In contrast, the relatively weaker ensemble-oscillator coupling strength ($g<g_{0}$) does not qualitatively alter our cooling protocol. However, it does affect the cooling rate, necessitating more probe spin projections to achieve the desired level of cooling. While these limitations make achieving full polarization with our protocol challenging, demonstrating partial polarization is expected to be feasible and would already constitute an exciting experiment.
Moreover, traditional methods, such as those based on resonant energy transfer, can initially be used to achieve a substantial degree of polarization in a spin ensemble. Our protocol can then be applied to further enhance the polarization to levels otherwise unattainable with energy transfer alone. This is because our protocol essentially projects out components with lower polarization, thereby increasing the success rate of obtaining desired probe measurements as the spin ensemble becomes more polarized. In the extreme case of an almost fully polarized spin ensemble, the success probability of a probe projection would be close to 1.
@@ V Discussion
In this study, we have developed a novel ground-state cooling protocol for a tripartite hybrid quantum system consisting of a spin ensemble, an oscillator, and a probe spin. Central to this protocol is the exploitation of the feedback provided by frequent measurements of the prob spin. Our analysis has shown that even with weak dispersive coupling between the spins and the oscillator, one can effectively cool both the ensemble and the oscillator to their ground states, either separately or simultaneously. Our analysis has also revealed a more complex relationship between the achievable polarization levels and the ensemble size. This relationship is further complicated by the detuning $\epsilon$, making it challenging to derive a straightforward scaling behavior for the required number of projections $M$ as a function of the ensemble size $N$. Establishing such scaling relations requires a more detailed analysis, which we leave it for future study.
Our cooling protocol offers several desirable features. First, it requires quantum coherence only in the probe, distinguishing it from resonant transfer schemes that demand coherence throughout the hybrid system. In addition, the periodic inversion pulses significantly enhance the probe spin’s protection, extending its coherence time well beyond current technological requirements. Furthermore, when applied to a macroscopic oscillator, this polarization technique can be viewed as a remote sensing scheme. This approach allows the sensing volume boundaries to extend beyond the limits imposed by phase accumulation schemes [12], offering potential advantages in various sensing applications.
@@
Numerical simulations are performed over a range of parameters, underscoring the feasibility of the cooling protocol and laying the groundwork for its experimental realization. For example, our results indicate that a spin ensemble consisting of $N=14$ spins could be cooled with a spin-mechanical coupling strength of only a few kHz at a mechanical oscillation frequency close to 1 MHz. Although such coupling strengths are within experimental reach, the cooling efficiency can be further improved by miniaturizing the mechanical oscillator [25]. In the context of solid-state defects in silicon carbide, coupling strengths can reach a few MHz [42], promising even more efficient cooling rates. Additionally, the use of magnons instead of NMOs could significantly enhance the spin-oscillator coupling, facilitating the experimental adoption of our protocol.
@@
A notable challenge of our cooling approach is its reliance on post-selection, which results in a decreasing success rate as the number of probe projections increases. Nevertheless, our protocol offers a significant advantage for cooling large spin ensembles that are traditionally difficult to handle and cannot be cooled by standard optical or energy transfer techniques. For spin ensembles that are amenable to conventional cooling methods, such as NV centers in diamond, our protocol expands the cooling toolbox by working in conjunction with other methods, thus mitigating the problem of exponentially decreasing success rate. In addition, our analysis of a specific NV cantilever system suggests the possibility of an exciting experimental demonstration of partial ensemble polarization using currently available technologies.
@@
This research highlights the significant impact that precise control of a single spin can have on the dynamics of a hybrid system, enabling energy transfer-like effects beyond the inherent capabilities of weak dispersive coupling. A similar idea has recently been demonstrated by an intriguing experiment: carefully designed inversion pulses on NV electronic spins can enable them to interact with extremely remote nuclear spins through paramagnetic spins (P1 centers) [13]. In addition to remote sensing, our results also provide valuable insights into the preparation of complex quantum states in a macroscopic object, such as the GKP states.
@@ Acknowledgement
The authors would like to thank Jixing Zhang, Xin Zhang, and Ruoming Peng for their valuable feedback on the manuscript. This work was supported by the German Federal Ministry of Education and Research (BMBF) through the project QECHQS with Grant No. 16KIS1590K, and through the Future Cluster QSens. Additional support was provided by the Baden-Württemberg Stiftung through the project SPOC with Grant agreement No. QT-6 and through the EU Horizon Project SPINUS (101135699). The Deutsche Forschungsgemeinschaft (DFG) has funded this work under GRK 2642, and under Project No. 384846402 as part of FOR 2724.
Appendix A Details of sketched experimental setup
The NV electronic spin is spin-1, which can be operated as a qubit with its $m_{s}=0,-1$ sublevels. This requires an external magnetic field aligned with its symmetry axis to separate the $m_{s}=\pm 1$ levels, which is typically around 400 Gauss for NV centers in diamond [44]. Such an arrangement negates the need to consider transverse spin-strain coupling due to its direct interaction between the $m_{s}=\pm 1$ levels [7, 20, 21], allowing us to safely ignore the $m_{s}=+1$ sublevel in this work. What remains is the longitudinal strain coupling acting on the qubit levels, with a number that can be up to 10 kHz for NV centers [48]. This number can be even higher for other solid-state defect centers [42].
In contrast, the spin ensemble to be cooled can be chosen more flexibly, since precise control of these spins is not a requirement for the current study. Potential targets include distant spins on other coupled cantilevers, spins at the clamped edge that are strain-coupled to the mechanical mode, or surface electron spins that are not well controlled. For example, if the ensemble spins are dark spins on the surface of the NMO [49] or other ferromagnetic particles [50], optical manipulation over them can be difficult, highlighting the need for other indirect polarization techniques. Note that this ensemble-oscillator interaction generated by the attached magnetic tip can be 1–2 orders of magnitude stronger than the spin-strain coupling between the probe and the oscillator [51].
Appendix B Effects of inhomogeneous coupling and imperfect readout
In this section, we briefly address the effects of other imperfections, such as inhomogeneous ensemble-oscillator coupling and imperfect probe readout. These factors can affect the cooling efficiency but do not impact the final achievable state.
.1 Inhomogeneous ensemble
The inhomogeneity in coupling between the ensemble spins and the oscillator primarily arises from the spatial distribution of the gradient generated by the external magnetic field or the strain environment. Ref. [53] demonstrates that for a spin ensemble located at a distance $h=10\,\text{nm}$ from a magnetic tip and distributed over a planar area of $1\,\mu\text{m}^{2}$, the resulting inhomogeneity can be approximately $g\pm g/10$. As the gradient based decreases with distance from the tip as, $1/r$, spins spread over larger distances are effectively uncoupled to the oscillator. Minimizing the intra-dipolar couplings to be smaller than $g_{0}$ among the spins, $N\approx 10$ spins. Despite the presence of inhomogeneous couplings, our scheme still enables cooling of the spin ensemble, albeit at a reduced speed, necessitating more successful probe projections.
As illustrated in the upper panel of Fig. 5, we sample an ensemble of four spins with randomly assigned coupling strengths to the oscillator, while keeping other parameters consistent with those in the main text. Compared to a uniformly coupled ensemble of the same size, both ensembles can be cooled to the ground state, although the inhomogeneous ensemble requires approximately 100 additional probe projections to achieve ground-state cooling.
.2 Imperfect probe readout
Imperfect readout of the probe spin (NV electronic spin considered in this work) often mainly occurs due to dark counts in the photodetector, which can mistakenly register incorrect projections as correct ones [31, 44]. However, these erroneous measurements primarily affect the cooling rate, as they typically result in restarting the cooling process. Importantly, they do not impact the final steady state of the oscillator and the spin ensemble.
In the bottom panel of Fig. 5, we illustrate the polarization of the spin ensemble reaching its maximum polarization sector, $N/2$, under conditions of imperfect projective measurements of the probe. We analyze a spin ensemble of four spins, each identically coupled to the oscillator. For the above simulations, the readout error is introduced as the probability that a projection to ’-’ is wrongly deemed as a successful probe projection into ‘+’ state. As the readout error rate increases, the cooling rate decreases, but the final steady state remains unaffected as shown in Fig. 5. Even with an average readout error rate of 0.25, the ensemble can still be cooled to the ground state relatively quickly. Given that the readout error rate for NV centers can be as low as 0.05 [31, 44], we believe that readout error will not pose a significant bottleneck in experimentally realizing our cooling scheme.
It is important to note that in the extreme case where both projections are equally probable (readout error rate of 0.5), the ensemble spins keep unpolarized, thereby reinforcing the validity of our measurement-based cooling scheme.
Figure 5: Polarizing spin ensemble in the presence of inhomogeneous ensemble-oscillator coupling and imperfect probe readout. The parameters involved in the simulation are $g/\omega=0.01$, $\epsilon/g=0.15$, and $n=100$ pulses per probe projection. (1) The upper panel shows the polarization of the spin ensemble reaching its maximum polarization sector, $N/2$, for both homogeneous and inhomogeneous couplings of the spins to the oscillator, with the same mean value. The inhomogeneous couplings are randomly sampled from a normal distribution with a mean value equal to $g$ and a standard deviation of $g/2$. In the homogeneous case, all four spins are identically coupled to the oscillator with strength $g$. (2) The bottom panel illustrates the polarization of the spin ensemble reaching its maximum polarization sector, $N/2$, under conditions of imperfect projective measurements of the probe. We consider a spin ensemble of four spins, each identically coupled to the oscillator. As the readout error increases, the cooling rate decreases, but the final steady state remains unaffected. In the extreme case where both projections are equally probable (error rate = 0.5), there is no cooling of the spin ensemble, confirming the validity of our measurement-based cooling scheme.
.3 Effects of spin-$T_{1}$ and oscillator $Q$
As the measurement based cooling scheme does not require coherences of the spin or the oscillator, the cooling fidelities are determined by their relative dissipation rates, $T_{1}$ for spins and $\gamma$ for the oscillator. In this work we have considered an ideal scenario i.e., infintie life time for the spins and the oscillator. Incorporating these effects will not be complex as no correlations are generated in the spin-mechanical system, and the probe measurements serve as an additional dissipation source for the oscillator and spin ensemble. The competition between the rates of their thermal equilibration and the measurement induced equilibration, results in a new steady state. For example in the presence of $T_{1}$, which tries to lower the maximal polarization $P^{z}_{N/2}$ towards $2/N$, at a rate $\gamma_{eq}$, and the measurement cooling increasing towards unity at a rate $\gamma_{c}$, the equilibrium polarization can be approximated using simple rate equations as
$P^{z}_{N/2}(M\rightarrow\infty)\sim\frac{\gamma_{c}}{\gamma_{c}+\gamma_{eq}}.$ (17)
For a typical range of $\gamma_{c}/\gamma_{eq}\geq 10$, the final polarization of the spin ensemble could always be larger than $90\%$. The dissipative effects due to the oscillator have a dominant effect on the polarization fidelity. As the measurement-based cooling is based on the memory/correlation time of the oscillator state, a rapidly relaxing oscillator ($\gamma_{osc}\sim g_{0}$) will not lead to any effective cooling of the spins.
Appendix C Simulation details